I need help with this question- what formula should I use and how to apply it?

Using the cash flows below, what is the NPV of all the cashflows using an interest rate of 10% ?

  • Now: - $120

  • After 1 year: + $20

  • After 2 years: + $90

  • After 3 years: - $10

  • After 4 years: + $30

  • After 5 years: + $30

  • After 6 years: + $30

  • After 7 years: + $40
Answer #1

Answer:

NPV = Sum of PV of all cashflow

PV =CF/(1+r)^n

where CF=cash flow

r=10%

n=year in which cash flow occur

For Year 2

n=2

CF=$90

PV=90/(1+10%)^2=$74.38

Year Cashflow PV
0 -120 -120.00
1 20 18.18
2 90 74.38
3 -10 -7.51
4 30 20.49
5 30 18.63
6 30 16.93
7 40 20.53
NPV 41.63